When 192.168.1.0/24 is divided into 8 subnets what is the size of each subnet?

16

64

32

none of the above
Answer #1

Answer

c)  32

Explanation:
-------------
new subnet mask is 27
so, size of each subnet = number of IP addresses per subnet block = 2^(32-27) = 2^5 = 32